OpenStudy (anonymous):

If h(x) = 2x + 3, find the value of x if h(x) = 9.

OpenStudy (lgbasallote):

if h (x) = 2x + 3 and h(x) is also = 9...then you can just say 2x + 3 = 9 right?

OpenStudy (anonymous):

u got it. thanks

OpenStudy (lgbasallote):

note that you need to solve for x

OpenStudy (mayankdevnani):

h(9)=2(9)+3 h(9)=18+3=21